# PyKirigami: An interactive Python simulator for kirigami metamaterials

### v2

@@ 4 Discussion
As illustrated in Fig. S1, a connection is defined by a pair of pivot points, such as $P$ on tile A and $Q$ on tile B. We provide PyBullet with the constant local coordinates of these pivots, $\mathbf{p}_{l}$ and $\mathbf{q}_{l}$. The physics engine then ensures that the world-space positions of these points, $\mathbf{p}_{w}$ and $\mathbf{q}_{w}$, satisfy the constraint equation
@@
S2.4 Interactive Control and Simulation State Management
@@
The Simulation class constructs all rigid bodies and constraints from files, configures damping/environment, and prepares data for runtime control. It returns a dictionary simulation_data for downstream controllers, summarized in Table S1.
Table S1: Contents of the simulation_data dictionary passed to controllers.
@@
For completeness and reproducibility, we summarize the default interactive controls used in PyKirigami’s GUI mode. These inputs allow manual actuation, dynamic boundary condition changes, camera manipulation, and on-demand export. Key bindings can be customized in the codebase; the mappings shown in Table S2 reflect the default configuration used in this paper.
Table S2: Interactive Control Scheme for the PyKirigami Simulator.
@@
S2.5 Physical Environment Construction: CLI Parameters and Defaults
This section documents the primary command-line interface (CLI) flags used to configure and run the simulations. The table below groups related parameters and lists their default values and usage. For simplicity and portability, input files can be organized into a single directory within the data/ folder and loaded using the --model flag. The CLI parameters and default values are given in Table S2.5:
Table S3: Command-Line Interface (CLI) parameters and default values. Parameter groups are distinguished by alternating background colors.
@@
Appendix S3 Example Configurations and Simulation Details
In the main text, we presented several 2D-to-2D, 2D-to-3D, and 3D-to-3D deployments obtained with PyKirigami. Here we provide additional snapshots and, crucially, the corresponding physical environment and solver settings used for each case so that readers can reproduce the results precisely. All CLI flags follow Table S2.5. All simulations were run with a fixed timestep $\Delta t=1/240$ s and $n_{\text{sub}}=20$ internal substeps unless otherwise noted.
S3.1 Example: 2D-to-2D Square-to-Circle Kirigami Model
The 2D-to-2D square-to-circle kirigami model example (Fig. S4; see also main text Fig. 4a) demonstrates a robust setup for a 2D radial expansion. The primary challenge is to enforce planar motion and prevent the structure from buckling out-of-plane. Our strategy combines hinge kinematics with specific environmental and geometric settings, driven by an automated radial force model.
Figure S4: Additional snapshots of the 2D-to-2D deployment of the square-to-circle kirigami model achieved by PyKirigami.
@@
Figure S5: Interactive workflow for the manual reconfiguration of the square-to-disk kirigami model. a, A central tile is pinned (Right-click) to create a stable reference frame. b, An outer tile is dragged (Left-click + drag) to begin the radial expansion. c, The structure is guided into an approximate “disk” shape. At this stage, small gaps often persist due to the implicit hinge forces. d, To close a gap, a tile is dragged into the correct position through Left-click + drag and the simulation is then paused (Space bar), and the tile is pinned (Right-click) while paused. This “pause-and-pin” technique is a powerful feature, allowing a constraint to be applied to a desired geometry before the physics solver can react. e, Step d is repeated for other misaligned tiles. f, All pins are removed while paused, leaving the structure in a stable and precisely “disk” state.
S3.2 Example: 2D-to-2D Square-to-Disk Kirigami Model (Interactive Workflow)
Fig. S5 demonstrates the fully interactive workflow for the 2D-to-2D compact reconfigurable square-to-disk model shown in main text Fig. 4b. While the main text analyzes this structure’s automated deployment, here we show how to achieve the same reconfiguration manually using the GUI controls detailed in Table S2. This process highlights the use of the simulator for hands-on exploration and state capture.
With the simulation running in a no-force mode, a user can manually reconfigure the structure from the “square” to the “disk” state. This is achieved by applying the interactive controls from Table S2 in a strategic sequence, as illustrated in Fig. S5.
@@
S3.3 Example: 2D-to-3D Square-to-Spherical-Cap Kirigami Model
The square-to-spherical-cap kirigami model example (Fig. S6; see also main text Fig. 4c) demonstrates a 2D-to-3D transformation from a flat sheet into a curved surface. This complex, out-of-plane motion is impossible to guide with simple radial forces and necessitates the use of the target-based actuation model for precise control. Here, additional snapshots showing the progression from a planar to a spherical configuration are provided. Also, one can notice that out-of-plane motion provides extra degrees of freedom, allowing the sheet to lift/twist to resolve incompatibilities which can be seen in [20].
Figure S6: Additional snapshots of the 2D-to-3D deployment of the square-to-spherical-cap kirigami model achieved by PyKirigami, driven by the target-based force model.
@@
S3.4 Example: 3D-to-3D Compact Reconfigurable Cylinder Kirigami Model
The 3D-to-3D compact reconfigurable cylinder kirigami model example (Fig. S7; see also main text Fig. 4d) involves the reconfiguration between two distinct 3D states, from a compact, folded cylinder to another compact, folded state. The target-based model is essential for guiding the structure through a collision-free volumetric transformation.
Figure S7: Additional snapshots of the reconfiguration process of the 3D-to-3D cylinder kirigami structure achieved by PyKirigami.
@@
S3.5 Example: 2D-to-2D Stampfli Quasicrystal Kirigami Model
Note that the above kirigami structures primarily consist of quadrilateral tiles, and the overall structures exhibit a highly regular topology. To further demonstrate the applicability of PyKirigami to other kirigami structures, in Fig. S8 we show the deployment of a 12-fold Stampfli quasicrystal kirigami structure from [16]. Here, note that the structure contains not only quadrilateral tiles but also triangular tiles, and the tile connection is also highly different from the ones shown above. It can be observed that PyKirigami is capable of producing a deployment simulation of this structure with automated radial expansion.
Figure S8: The simulated deployment of the 12-fold Stampfli quasicrystal kirigami structure achieved by PyKirigami with automated radial expansion, stabilized by environmental forces.
@@
S3.6 Example: 2D-to-2D Compact Reconfigurable Fan Kirigami Model
To further demonstrate the versatility of PyKirigami, we present another example focusing on its interactive reconfiguration functionality. Fig. S9 shows the compact reconfigurable fan kirigami model, which has two different closed and compact contracted states. Users can apply the mentioned interactive operations to deploy it under the same physical environment as the square-to-disk model as shown in Section S3.2.
Figure S9: The simulated deployment of the compact reconfigurable fan kirigami model produced by PyKirigami with interactive shape morphing.
Figure S10: The simulated deployment of the cube-to-sphere kirigami model produced by PyKirigami. In this simulation, large, coordinated out-of-plane rotations and precise positioning are applied to control the deployment effect.
S3.7 Example: 3D-to-3D cube-to-sphere Kirigami Model
Finally, we present a challenging 3D-to-3D transformation example using the cube-to-sphere kirigami model (Fig. S10; see also main text Fig. 1b). For this model, the six faces of the cube are deployed into a spherical approximation throughout the deployment process, which requires large, coordinated out-of-plane rotations and precise positioning to ensure the edges meet correctly without self-intersection. Such a transformation is infeasible with simple or manual forces and serves as a powerful demonstration of the target-based actuation model’s capabilities.
@@
Appendix S4 Detailed Analysis of Simulation Parameters and Residual Energy
S4.1 Residual energy on a logarithmic scale
Fig. S11 and Fig. S12 show the log-scale counterparts of the energy plots in main text Fig. 4 and Fig. 5b. The small oscillations seen near the compact shape reflect a dynamic equilibrium: the constraint solver performs continuous, small corrective work at the joints, while damping and friction dissipate energy. On a linear scale, this residual is several orders of magnitude below the deployment peak and appears negligible; the log view simply magnifies it.
Figure S11: Log-scale energy plot for the passive dynamics in the square-to-disk deployment simulation. The residual “floor” is the balance between tiny constraint corrections and dissipation.
Figure S12: Log-scale energy plot for the actively controlled deployment ($k_{s}=100$) of the square-to-disk kirigami model. Residual micro‑oscillations persist but are orders of magnitude below the peak.
S4.2 Friction Budget and the Underactuation Threshold
In PyBullet, each rigid body is assigned a lateralFriction coefficient. The effective coefficient of friction, $\mu_{\mathrm{eff}}$, at a contact point between two bodies is determined by the product of their individual coefficients:
$\mu_{\mathrm{eff}}=\mu_{\text{tile}}\cdot\mu_{\text{ground}}.$ (5)
With the default values used in our experiments, $\mu_{\text{tile}}=\mu_{\text{ground}}=0.5$, which results in an effective friction coefficient of $\mu_{\mathrm{eff}}=0.25$.
On a horizontal plane, the maximum tangential static friction force, $F_{\mathrm{fric}}^{\max}$, that can be exerted on a tile is bounded by the normal force $F_{n}$:
$F_{\mathrm{fric}}^{\max}=\mu_{\mathrm{eff}}\,\|F_{n}\|.$ (6)
For a tile of mass $m=1\,\mathrm{kg}$ under a gravitational acceleration of magnitude $g=-200\,\mathrm{m/s^{2}}$, the normal force per tile is $F_{n}=mg=-200\,\mathrm{N}$. The maximum static friction force per tile is therefore
$F_{\mathrm{fric}}^{\max}=\|\mu_{\mathrm{eff}}\cdot mg\|=50\,\mathrm{N}.$ (7)
Static friction reacts to oppose any applied tangential force up to this limit. Because multiple tiles are in contact with the ground, the total frictional resistance is the sum over all contacting tiles. This friction budget quantitatively explains the “underactuated” outcome in the main text (Fig. 5a): with a low spring stiffness of $k_{s}=50$, the actuation forces are insufficient to overcome the combined resistance from the structure’s intrinsic potential energy landscape and the cumulative ground friction. As a result, the structure stalls in an intermediate state.
Appendix S5 Video Captions

### v3

@@ B.3 Creating Inter-Tile Constraints
As illustrated in Fig. 7, a connection is defined by a pair of pivot points, such as $P$ on tile A and $Q$ on tile B. We provide PyBullet with the constant local coordinates of these pivots, $\mathbf{p}_{l}$ and $\mathbf{q}_{l}$. The physics engine then ensures that the world-space positions of these points, $\mathbf{p}_{w}$ and $\mathbf{q}_{w}$, satisfy the constraint equation
@@
B.4 Interactive Control and Simulation State Management
@@
The Simulation class constructs all rigid bodies and constraints from files, configures damping/environment, and prepares data for runtime control. It returns a dictionary simulation_data for downstream controllers, summarized in Table 1.
Table 1: Contents of the simulation_data dictionary passed to controllers.
@@
For completeness and reproducibility, we summarize the default interactive controls used in PyKirigami’s GUI mode. These inputs allow manual actuation, dynamic boundary condition changes, camera manipulation, and on-demand export. Key bindings can be customized in the codebase; the mappings shown in Table 2 reflect the default configuration used in this paper.
Table 2: Interactive Control Scheme for the PyKirigami Simulator.
@@
B.5 Physical Environment Construction: CLI Parameters and Defaults
This section documents the primary command-line interface (CLI) flags used to configure and run the simulations. The table below groups related parameters and lists their default values and usage. For simplicity and portability, input files can be organized into a single directory within the data/ folder and loaded using the --model flag. The CLI parameters and default values are given in Table B.5:
Table 3: Command-Line Interface (CLI) parameters and default values. Parameter groups are distinguished by alternating background colors.
@@ B.5 Physical Environment Construction: CLI Parameters and Defaults
Appendix C Example Configurations and Simulation Details
In the main text, we presented several 2D-to-2D, 2D-to-3D, and 3D-to-3D deployments obtained with PyKirigami. Here we provide additional snapshots and, crucially, the corresponding physical environment and solver settings used for each case so that readers can reproduce the results precisely. All CLI flags follow Table B.5. All simulations were run with a fixed timestep $\Delta t=1/240$ s and $n_{\text{sub}}=20$ internal substeps unless otherwise noted.
C.1 Example: 2D-to-2D Square-to-Circle Kirigami Model
The 2D-to-2D square-to-circle kirigami model example (Fig. 10; see also main text Fig. 3a) demonstrates a robust setup for a 2D radial expansion. The primary challenge is to enforce planar motion and prevent the structure from buckling out-of-plane. Our strategy combines hinge kinematics with specific environmental and geometric settings, driven by an automated radial force model.
Figure 10: Additional snapshots of the 2D-to-2D deployment of the square-to-circle kirigami model achieved by PyKirigami.
@@ C.1 Example: 2D-to-2D Square-to-Circle Kirigami Model
Figure 11: Interactive workflow for the manual reconfiguration of the square-to-disk kirigami model. a, A central tile is pinned (Right-click) to create a stable reference frame. b, An outer tile is dragged (Left-click + drag) to begin the radial expansion. c, The structure is guided into an approximate “disk” shape. At this stage, small gaps often persist due to the implicit hinge forces. d, To close a gap, a tile is dragged into the correct position through Left-click + drag and the simulation is then paused (Space bar), and the tile is pinned (Right-click) while paused. This “pause-and-pin” technique is a powerful feature, allowing a constraint to be applied to a desired geometry before the physics solver can react. e, Step d is repeated for other misaligned tiles. f, All pins are removed while paused, leaving the structure in a stable and precisely “disk” state.
C.2 Example: 2D-to-2D Square-to-Disk Kirigami Model (Interactive Workflow)
Fig. 11 demonstrates the fully interactive workflow for the 2D-to-2D compact reconfigurable square-to-disk model shown in main text Fig. 3b. While the main text analyzes this structure’s automated deployment, here we show how to achieve the same reconfiguration manually using the GUI controls detailed in Table 2. This process highlights the use of the simulator for hands-on exploration and state capture.
With the simulation running in a no-force mode, a user can manually reconfigure the structure from the “square” to the “disk” state. This is achieved by applying the interactive controls from Table 2 in a strategic sequence, as illustrated in Fig. 11.
@@ C.2 Example: 2D-to-2D Square-to-Disk Kirigami Model (Interactive Workflow)
C.3 Example: 2D-to-3D Square-to-Spherical-Cap Kirigami Model
The square-to-spherical-cap kirigami model example (Fig. 12; see also main text Fig. 3c) demonstrates a 2D-to-3D transformation from a flat sheet into a curved surface. This complex, out-of-plane motion is impossible to guide with simple radial forces and necessitates the use of the target-based actuation model for precise control. Here, additional snapshots showing the progression from a planar to a spherical configuration are provided. Also, one can notice that out-of-plane motion provides extra degrees of freedom, allowing the sheet to lift/twist to resolve incompatibilities which can be seen in [20].
Figure 12: Additional snapshots of the 2D-to-3D deployment of the square-to-spherical-cap kirigami model achieved by PyKirigami, driven by the target-based force model.
@@ C.3 Example: 2D-to-3D Square-to-Spherical-Cap Kirigami Model
C.4 Example: 3D-to-3D Compact Reconfigurable Cylinder Kirigami Model
The 3D-to-3D compact reconfigurable cylinder kirigami model example (Fig. 13; see also main text Fig. 3d) involves the reconfiguration between two distinct 3D states, from a compact, folded cylinder to another compact, folded state. The target-based model is essential for guiding the structure through a collision-free volumetric transformation.
Figure 13: Additional snapshots of the reconfiguration process of the 3D-to-3D cylinder kirigami structure achieved by PyKirigami.
@@ C.4 Example: 3D-to-3D Compact Reconfigurable Cylinder Kirigami Model
C.5 Example: 2D-to-2D Stampfli Quasicrystal Kirigami Model
Note that the above kirigami structures primarily consist of quadrilateral tiles, and the overall structures exhibit a highly regular topology. To further demonstrate the applicability of PyKirigami to other kirigami structures, in Fig. 14 we show the deployment of a 12-fold Stampfli quasicrystal kirigami structure from [16]. Here, note that the structure contains not only quadrilateral tiles but also triangular tiles, and the tile connection is also highly different from the ones shown above. It can be observed that PyKirigami is capable of producing a deployment simulation of this structure with automated radial expansion.
Figure 14: The simulated deployment of the 12-fold Stampfli quasicrystal kirigami structure achieved by PyKirigami with automated radial expansion, stabilized by environmental forces.
@@ C.5 Example: 2D-to-2D Stampfli Quasicrystal Kirigami Model
C.6 Example: 2D-to-2D Compact Reconfigurable Fan Kirigami Model
To further demonstrate the versatility of PyKirigami, we present another example focusing on its interactive reconfiguration functionality. Fig. 15 shows the compact reconfigurable fan kirigami model, which has two different closed and compact contracted states. Users can apply the mentioned interactive operations to deploy it under the same physical environment as the square-to-disk model as shown in C.2.
Figure 15: The simulated deployment of the compact reconfigurable fan kirigami model produced by PyKirigami with interactive shape morphing.
Figure 16: The simulated deployment of the cube-to-sphere kirigami model produced by PyKirigami. In this simulation, large, coordinated out-of-plane rotations and precise positioning are applied to control the deployment effect.
C.7 Example: 3D-to-3D cube-to-sphere Kirigami Model
Finally, we present a challenging 3D-to-3D transformation example using the cube-to-sphere kirigami model (Fig. 16; see also main text Fig. 1b). For this model, the six faces of the cube are deployed into a spherical approximation throughout the deployment process, which requires large, coordinated out-of-plane rotations and precise positioning to ensure the edges meet correctly without self-intersection. Such a transformation is infeasible with simple or manual forces and serves as a powerful demonstration of the target-based actuation model’s capabilities.
@@ C.7 Example: 3D-to-3D cube-to-sphere Kirigami Model
Appendix D Analytic Kinematic Models for Quantitative Validation
Figure 17: The definition of the deployment angle in the benchmark cases.a. The rectangular tessellation kirigami structure. The dashed line shows a unit cell of the tessellation, in which the deployment angle $\theta$ can be defined. b, The rigidly-deployable square-to-disk kirigami structure from [25], in which the deployment can be characterized by a single deployment angle $\theta$.
To validate the numerical accuracy of PyKirigami, we derive the exact kinematic equations for the two benchmark kirigami structures used in main text. Both systems function as single-degree-of-freedom (1-DOF) mechanisms, where the global configuration is uniquely determined by the deployment angle $\theta$.
D.1 Rectangular Tessellation
The rectangular tessellation is frequently used as a base case in kirigami design, and its analytical vertex position along with deployment angle $\theta$ can be easily derived. As shown in Fig. 17a, the unit cell consisting of four rectangles has the size $L\times W$ where (LW ) = 2 (sin(η) cos(η)cos(η) sin(η) )(lw ),  η= θ/2, assuming that the small rectangles have size $l\times w$.
D.2 Square-to-Disk Kirigami Structure
The compact reconfigurable square-to-disk kirigami structure exhibits a more complex, non-linear kinematic path. The radial position $R$ of boundary vertices is coupled to the deployment angle $\theta$ (Fig. 17b). The position of all inner vertices $\mathbf{v}_{k}$ in the assembly can be derived via a matrix multiplication v_k = M V, where the matrix $M$ is related to the deployment angle and $V$ are the seed vertices. More details can be found in [25].
Appendix E Video Captions